In a coffee cup calorimeter, 5.86 g of SO3 (g) (mm SO3=80.06 g/mol) is reacted with 100.00g of water at 25.0 Celsius. Assuming the specific heat for the solution is 4.184 J/gC and the heat of reaction for 1.00 mole of SO3 (g) is -241.8 kJ/mol, what is the final temperature of the given reaction solution?
Answer #1

moles of SO3 = 5.86/ 80.06 = 0.0732

Q    = 241.8 x 0.0732 = 17.7 kJ

Q =17699 J

total mass = 105.86 g

Q = m cpdT

17699 = 105.86 x 4.184 x dT

dT = 39.96

T2 = 39.96 + 25

T2 = 65.0 oC

final temperature = 65.0oC